# Can Community Notes Replace Professional Fact-Checkers?

### v1

@@ Acknowledgements
Appendix A Additional Material
@@
Figure 7: The categories of links used by Community notes’ authors as a source, filtering for notes rated as “helpful”.
Figure 8: The categories of links used by Community notes’ authors as a source, filtering for notes rated as “not helpful”.
@@
Domain Category Domain Category
x.com social media thehill.com news
twitter.com social media amp.theguardian.com news
youtube.com social media whitehouse.gov government
youtu.be social media news.sky.com news
un.org organisation merriam-webster.com reference
u.today news techarp.com news
t.co social media cbc.ca news
snopes.com fact checking politifact.com fact checking
en.m.wikipedia.org reference pbs.org commercial
en.wikipedia.org reference telegraph.co.uk news
google.com search engine businessinsider.com news
instagram.com social media time.com news
britannica.com reference justice.gov government
reuters.com news cnbc.com news
bbc.co.uk news wsj.com news
apnews.com news sciencedirect.com academic
bbc.com news msn.com news
nytimes.com news statista.com reference
theguardian.com news business.x.com commercial
vice.com news amp.cnn.com news
usatoday.com news congress.gov government
factcheck.org fact checking factcheck.afp.com fact checking
cnn.com news yahoo.com search engine
washingtonpost.com news timesofindia.indiatimes.com news
ncbi.nlm.nih.gov academic thelancet.com academic
nbcnews.com news hrw.org organisation
help.twitter.com reference healthfeedback.org fact checking
cdc.gov government fda.gov government
npr.org news m.youtube.com social media
forbes.com news law.cornell.edu academic
newsweek.com news medium.com blog post
fullfact.org fact checking healthfeedback.org fact checking
dailymail.co.uk news who.int organisation
cbsnews.com news haaretz.com news
web3antivirus.io database axios.com news
timesofisrael.com news mayoclinic.org commercial
help.x.com reference nejm.org academic
nypost.com news scienceexchange.caltech.edu academic
aljazeera.com news indiatoday.in news
reddit.com social media bloomberg.com news
independent.co.uk news pewresearch.org academic
usgs.gov academic jamanetwork.com academic
abcnews.go.com news leadstories.com news
nature.com academic dictionary.cambridge.org reference
gov.uk government jpost.com news
web.archive.org database archive.ph database
foxnews.com news healthline.com commercial
tiktok.com social media abc.net.au news
edition.cnn.com news france24.com news
Table 4: List of top 100 most common domains found in the community notes dataset, and their categorization.
@@
Fig. 7 The categories of links used by Community notes’ authors as a source, filtering for notes rated as “helpful”.
Fig. 8 The categories of links used by Community notes’ authors as a source, filtering for notes rated as “not helpful”.
@@
Tab. 3 List of top 100 most common domains found in the community notes dataset, and their categorization.
@@ Appendix B Reproducibility
B.1 Manual Annotation Setup

### v2

@@ 5 Conclusion
Community notes have been proposed as a replacement for professional fact-checkers and a salve to some of the issues encountered by fact-checking. However, our findings support the view that neither community notes nor professional fact-checkers alone are sufficient to combat the spread of misinformation on social media. Rather, a combination of these two strategies could prove a much more effective approach to addressing the full range of false content shared online, for example, by leveraging community notes to identify new checkworthy claims for professional fact-checkers, or relying on fact-checkers’ expertise to resolve disputed unpublished notes (see Augenstein et al. (2025) for further recommendations in this vein). In particular, as discussed above, professional fact-checking organisations are especially vital for verifying claims related to broader narratives and conspiracy theories. Moreover, professional fact-checkers remain the only viable strategy currently available for addressing partisan issues, where community notes fall short. Finally, we highlight the potential for incorporating automated, human-in-the-loop fact-checking models to assist professional and community fact-checkers alike in reckoning with vast amounts of both human- and machine-generated content.
@@ Acknowledgements
Appendix A Extended Background
A.1 Impact of Community Notes on misinformation spread
People shown community notes alongside misleading social media posts were more accurate in identifying misleading posts, and the notes were judged to be more trustworthy than context-free misinformation flags (e.g., "Checked by fact-checkers" or "Checked by other social media users"), regardless of (US-centric) political beliefs Drolsbach et al. (2024). People shown either community notes or related news article suggestions were both less likely to believe and report misleading information compared to a control group: community notes were more effective in reducing belief and sharing intention for positive rumours, while articles were more effective for negative rumours Kankham and Hou (2024). On the other hand, displaying community notes leads users to post more negative and angry replies to misleading posts Chuai et al. (2024b), while crowd workers are also prone to cognitive biases, such as overestimating a statement’s truthfulness the more they liked its claimant, and general overconfidence in their ability to ascertain truthful statements Draws et al. (2022).
A.2 Professional fact-checking and community note practices
Although fact-checks and community notes share similarities in how they address misleading claims, they also differ in key elements of practice and communication (Kankham and Hou, 2024). Fact-checking typically involves the analysis and verification of public claims, and in addition to verifying claims, in recent years many fact-checking organisations have also assumed a wider role in combating misinformation spread, e.g., long-term investigative journalism projects, media literacy programs (Juneja and Mitra, 2022). Professional fact-checkers signatory to the International Fact-Checking Network follow a rigorous set of principles and transparency commitments [^11] https://www.ifcncodeofprinciples.poynter.org/the-commitments and a structured workflow: (i) claim selection; (ii) collecting evidence; (iii) deciding on a verdict; and (iv) writing the fact-checking article (Graves, 2017; Micallef et al., 2022; Warren et al., 2025). In contrast, any platform user can contribute to community notes under anonymity, and the rating approach relies on the ‘wisdom of crowds’, with little oversight or transparency regarding biases of note-writers. Community note writers and fact-checkers tend to target similar topics (e.g., health and politics) (Saeed et al., 2022). Fact-checkers must rely on credible sources and evidence to convince the reader, while note writers often disagree with fact-checkers on what constitutes a reliable source, particularly on political ideological grounds Saeed et al. (2022). The verdicts reached in notes tend to agree with those of fact-checkers — however, due to political polarisation (Yasseri and Menczer, 2023), community notes on contentious political issues rarely reach a consensus (Saeed et al., 2022). Fact-checking articles, which are subject to multiple rounds of editorial scrutiny, are more formal and standardised in style than community notes, which vary considerably and can employ a range of persuasion techniques, such as appeals to emotion or other logical fallacies (Kankham and Hou, 2024). Moreover, community notes typically serve as direct rebuttals to misleading posts, while fact-checking articles may address a more general claim than is expressed in a specific post. Finally, fact-checking articles are a one-way exchange, while community notes represent a more horizontal and interactive dialogue between writer and recipient of the fact-check (Kankham and Hou, 2024). Prior work has found that collaborative fact-checking, a distinct approach to community notes for its Wikipedia-style approach that allows users to edit, as well as up and downvote user-written posts (Haime, 2022), can produce fact-checks with comparable speed, reliability, objectivity, clarity and persuasiveness to those written by professional fact-checks (Zhao and Naaman, 2023). However, laypeople’s work is expedited by existing fact-checking articles, and amateurs tend to defer to professional fact-checkers for topics requiring specific expertise, such as medical claims (Zhao and Naaman, 2023). Our work builds on current understanding of the relationship between professional fact-checking and community moderation by examining the extent to which community note writers deploy the work of fact-checkers in their notes.
Appendix B Additional Material
@@
Domain Category Domain Category
x.com social media thehill.com news
twitter.com social media amp.theguardian.com news
youtube.com social media whitehouse.gov government
youtu.be social media news.sky.com news
un.org organisation merriam-webster.com reference
u.today news techarp.com news
t.co social media cbc.ca news
snopes.com fact checking politifact.com fact checking
en.m.wikipedia.org reference pbs.org commercial
en.wikipedia.org reference telegraph.co.uk news
google.com search engine businessinsider.com news
instagram.com social media time.com news
britannica.com reference justice.gov government
reuters.com news cnbc.com news
bbc.co.uk news wsj.com news
apnews.com news sciencedirect.com academic
bbc.com news msn.com news
nytimes.com news statista.com reference
theguardian.com news business.x.com commercial
vice.com news amp.cnn.com news
usatoday.com news congress.gov government
factcheck.org fact checking factcheck.afp.com fact checking
cnn.com news yahoo.com search engine
washingtonpost.com news timesofindia.indiatimes.com news
ncbi.nlm.nih.gov academic thelancet.com academic
nbcnews.com news hrw.org organisation
help.twitter.com reference healthfeedback.org fact checking
cdc.gov government fda.gov government
npr.org news m.youtube.com social media
forbes.com news law.cornell.edu academic
newsweek.com news medium.com blog post
fullfact.org fact checking healthfeedback.org fact checking
dailymail.co.uk news who.int organisation
cbsnews.com news haaretz.com news
web3antivirus.io database axios.com news
timesofisrael.com news mayoclinic.org commercial
help.x.com reference nejm.org academic
nypost.com news scienceexchange.caltech.edu academic
aljazeera.com news indiatoday.in news
reddit.com social media bloomberg.com news
independent.co.uk news pewresearch.org academic
usgs.gov academic jamanetwork.com academic
abcnews.go.com news leadstories.com news
nature.com academic dictionary.cambridge.org reference
gov.uk government jpost.com news
web.archive.org database archive.ph database
foxnews.com news healthline.com commercial
tiktok.com social media abc.net.au news
edition.cnn.com news france24.com news
Table 2: List of top 100 most common domains found in the community notes dataset, and their categorisation.
@@
Tweet Note
misleadingUnverifiedClaimAsFact
misleadingOutdatedInformation
misleadingFactualError
misleadingSatire
Fact Checking source
@cnviolations I swear community notes are the only good thing Elon added since he bought Twitter. Community notes was first launched under former Twitter CEO Jack Dorsey in 2021 under the name of ‘Birdwatch’. The only thing Elon Musk did was that he renamed the feature to community notes. https://blog.twitter.com/en_us/topics/product/2021/introducing-birdwatch-a-community-based-approach-to-misinformation https://www.reuters.com/article/factcheck-elon-birdwatch-idUSL1N31Z2VG/
Thailand will become the first country to make the contract null and void, meaning that Pfizer will become responsible for all vaccine injuries … Thailand has no plans to void its Pfizer COVID vaccine contract, an official with the country’s National Vaccine Institute said. Thailand’s Department of Disease Control also rejected the claims as ‘fake news.’ … https://apnews.com/article/fact-check-covid-vaccine-pfizer-thailand-203948163859
@@
Tab. 2 List of top 100 most common domains found in the community notes dataset, and their categorisation.
@@
Tab. 4 A sample of tweets, notes, and their community annotations, as well as whether the note contains a fact-checking link.
@@ Appendix C Reproducibility
C.1 Manual Annotation Setup